# TechCoach: Towards Technical Keypoint-Aware Descriptive Action Coaching

### v1

@@ 5.3 Main Results
Table 5: Ablation studies on the impact of Keypoint Features Augmentation. K.P.Text: Keypoint text features. K.P.Type: keypoint-type prompts. Quality: quality-related prompts.
Variants Score Commentary
@@
Figure 5: Qualitative results on the generated coaching commentary. Our TechCoach generates precise and detailed coaching commentary on what is done well and what can be improved from the given action videos. Correctly matched parts are highlighted in colors.
5.4 Ablation Studies
- Is keypoint-level alignment a necessary? We study this question by gradually ablating the training losses illustrated in Eq. 3 from our full model. The results are shown in Tab. 3. By comparing Row 3 with Row 1-2, it can be clearly observed that the model achieves unstable performance when simply combining the $L_{MSE}$ and $L_{MTM}$. This may caused by the gap between the score regression and commentary generation tasks. However, when further combining the keypoint-level alignment loss (Row 4), we can see stable performance improvement on all the compared metrics. Hence, we can conclude that keypoint-level alignment is necessary for the trade-off of score regression and coaching commentary generation.
- Does keypoint-level alignment equal to binary classification? The main target of the keypoint-level alignment (KP-Align) is to ensure the Keypoint-related Quality Embeddings (see Fig. 3 and Sec. 4.4 for more details) carry the quality-related information. To achieve this, a simple solution is to perform binary classifications over the Keypoint-related Quality Embeddings (i.e., classify whether the execution shows strengths or weaknesses on each keypoint), which we note as KP-CLS. To go deeper, we conduct a study on the influence of such an alternative solution. As shown in Tab. 4, by comparing Row 2 with Row 1, we observe such a solution strengthens the quality perception of the model, resulting in performance improvement on score regression metrics. However, without explicit alignment with the keypoint-level coaching commentary, Bi-CLS falls short in understanding deeper relationships between the action execution and keypoints, resulting in a clear performance gap between our keypoint-level alignment, especially on commentary generation.
- Impacts of the Keypoint Query Augmentation. As discussed in Sec. 4.4, we construct keypoint queries by augmenting the keypoint text features (K.P.Text) with keypoint-type prompts (K.P.Type) and quality-related prompts (Quality). As shown in Tab. 5, we gradually ablate the components from the augmentation process, and the performance deterioration shows the contribution of each component.
@@ 5.4 Ablation Studies
- Impacts of videos from various views. As discussed in Sec. 5.2, we train TechCoach with both Egocentric and Exocentric videos. Here we conduct an experiment on the influence of the training videos from different views. As shown in Tab. 6, Ego-Only and Exo-Only variants showcase comparable performance. By combining videos from different views, TechCoach achieves stronger performance, which shows that egocentric and exocentric viewpoints are complementary for action coaching.
5.5 Qualitative Results
To further illustrate the effectiveness of the proposed TechCoach, we provide qualitative results on the generated coaching commentary. As shown in Fig. 5, TechCoach is able to understand various technical keypoints of different actions and provide detailed commentary on what is done well (e.g., “the extension of the shooting arm and wrist action were also commendable”) and what can be improved (e.g., “improvements are needed in core engagement to prevent being pulled from the wall”) from the given action videos. More qualitative results are shown in the Appendix.
6 Conclusion
In this paper, we investigate Descriptive Action Coaching, a novel task that focuses on providing coaching feedback on what is done well and what can be improved from an action execution. To achieve this, we propose TechCoach, a new framework that incorporates keypoint-level reasoning into the action coaching process, just as human coaches do. To support this task, we devise an automated pipeline for constructing the EE4D-DAC dataset which features clean and detailed coaching commentary on both keypoint and instance levels. Extensive experiments highlight the strong performance of the proposed TechCoach and demonstrate the effectiveness and the necessity of the proposed keypoint-level reasoning. We expect the proposed new task, dataset, and method provide a new perspective for advancing current AQA to a more explainable and practical scenario.
Appendix A Appendix
In this appendix, we provide more details about the evaluation metrics, implementations, and qualitative results in Appendix B. Subsequently, more details about the EE4D-DAC dataset are introduced in Appendix C, including examples of general technical keypoints, hierarchical coaching commentary, prompts used for LLM-based annotation collection, and other details. Finally, we discuss the limitations and potential directions for future work in Appendix D.
Appendix B More Details about Experiments
B.1 More Details about the Evaluation Metrics
Spearman’s rank correlation coefficient ($\rho$) [38] is the most widely used metric to evaluate models in the field of AQA. It measures the correlation between the predicted score series and the ground-truth score series, which is defined as follows:
$\rho=\frac{\sum^{N}_{i=1}(x_{i}-\bar{x})(y_{i}-\bar{y})}{\sqrt{\sum^{N}_{i=1}(x_{i}-\bar{x})^{2}\sum^{N}_{i=1}(y_{i}-\bar{y})^{2}}},$ (4)
where $x$ and $y$ are the rankings of the two series.
Relative-L2 Distance (RL2) [50] evaluates the numerical differences between the predicted and ground-truth scores, which can be written as:
$RL2=\frac{1}{N}\sum^{N}_{n=1}(\frac{|s_{n}-\hat{s}_{n}|}{s_{max}-s_{min}})^{2},$ (5)
where $s_{n}$ and $\hat{s}_{n}$ indicate the ground-truth and predicted scores, respectively. $s_{max}$ and $s_{min}$ indicate the highest and lowest scores for an action instance.
BLEU [25] quantifies the overlap between the generated sentences and the reference sentences by comparing n-grams (contiguous word sequences) shared between them. In this work, we compute the 1-gram-based BLEU score to evaluate the models.
METEOR [4] improves upon BLEU by introducing flexible word matching (e.g., synonyms and stemming) to better capture meaning beyond exact matches. Additionally, it adopts a recall-oriented approach to ensure all reference words are considered, and incorporates word order penalties to evaluate fluency and coherence.
CIDEr [37] is originally proposed for image captioning task. It measures the similarity between generated and reference texts using n-gram overlap weighted by TF-IDF. This ensures important but less common phrases are emphasized, rewarding consensus among references while penalizing overly generic outputs.
BERTScore [56] evaluates natural language generation by comparing the semantic similarity between generated and reference texts using contextual embeddings from pre-trained BERT-like models. Unlike traditional metrics that rely on n-gram overlap, BERTScore computes token-level similarity based on meaning, capturing nuances like synonyms and paraphrasing. In our work, we use the DeBERTa-xlarge [11] to extract token-level embeddings as it shows great correlation with human evaluation.
Figure 6: The detailed prompts for LLM-based evaluation. Zoom in for best view.
Figure 7: An overview of the Spatial-Temporal Context Enhancer. Here the $\oplus$ indicates the element-level adding operation.
GPT-M and GPT-Q are the LLM-based metrics proposed to evaluate whether the generated commentary shares the same technical details and the same coaching opinions as the ground-truth commentary. We employ the GPT-4o-mini to determine the number of both-mentioned technical details (i.e., keypoints) and the shared coaching opinions. Detailed prompts are provided in Fig. 6. After that, the GPT-M score is calculated as the averaged proportion of the both-mentioned technical details to those mentioned by ground truth across all testing instances. The GPT-Q score is calculated as the averaged proportion of the technical details sharing the same coaching opinions to the both-mentioned technical details.
B.2 More Implementation Details
Spatial-Temporal Context Enhancer. As discussed in Sec. 4.3 of the main paper, inspired by TimeSformer [5], we adopt a Spatial-Temporal Context Enhancer to enhance the spatial-temporal context across the clip-level features. An overview of this module is shown in Fig. 7. Specifically, we use a linear projector to map the clip-level features $F_{clip}$ to the hidden dimension $D$, and feed them into $N_{st}$ Transformer blocks that include spatial- and temporal-attention layers along with the FFN layers. In our implementation, the $N_{st}$ is set to 4.
Other details of TechCaoch. The visual backbone and text encoder are initialized by the pretrained weights and frozen during training. For the visual backbone, we use the pretrained weights of InternVideo2-stage3 [41]. Following [17, 54], we employ AdamW [18] as the optimizer, and utilize a learning rate warm-up during the early 10% training steps followed by linear decay. $\lambda_{1}$ and $\lambda_{2}$ are set to 0.1 and 0.3, respectively. Due to the characteristics of EgoExo4D [10] and the annotation pipeline of EE4D-DAC, it is not possible to ensure each instance includes rounded commentary on every possible technical keypoints (e.g., in some cases the coaching commentary would not mention the execution about the head & eyes). Therefore, we make some adjustments on the prompts or instructions for the compared models to guide the models to focus on those technical keypoints mentioned in the ground-truth Commentary. For TechCoach, we employ zero-masking on the keypoint-related quality embeddings corresponded to a specific keypoint only if both the strength and weakness on this keypoint are unmentioned in the ground truth. Note that such a design does NOT leak the action quality information, and similar adjustments are applied on the compared models [54, 16, 14, 41, 6] for fair comparisons.
Details about USDL [35]. USDL proposes to convert the score regression problem to a distribution prediction problem. We primarily follow the official implementation. We make adjustments on the output dimension to align with the score range in our experiments.
Details about CoRe [50] and TPT [3]. These two methods adopt the Contrastive Regression paradigm to generate action quality scores by predicting the score differences between the query instance and multiple exemplar instances. Our experiments are conducted based on the official implementations of these methods. The exemplars are randomly selected from the training instances with the same action task (e.g., Reverse Layup) as the query instance. The number of selected exemplars is set to 10.
Details about SwinBERT [17]. SwinBERT proposes a general framework for video captioning. In addition to the modification of the visual backbone, we also modified the length of masking tokens during training to ensure consistency with our TechCoach. Additionally, following [54], we add an extra full-connected layer to regress the score.
Details about PGMI [54]. PGMI is a prompt-guided framework for Multi-Task Learning AQA paradigm. Our experiments on PGMI are conducted using the official implementation. We adjust the number of score groups to 10 as the score range in our experiments is in [0, 10]. Besides, we modify the input templates for the text generator to include the score information and the technical dimensions that should be focused on. Specifically, the template is adjusted as follows:
The player is practicing <Action_Task>, and gets a score of <Predicted_Score>. Pay attention to the execution of <Technical_Dimensions>”,
where <Action_Task> is the action task, <Predicted_Score> indicates the predicted score by the model, and <Technical_Dimensions> indicates the a list of technical dimensions (e.g., head & eyes, arms & elbows, and human-object interaction) whose corresponded technical keypoints are mentioned in the ground truth.
Details about the Compared MLLMs [16, 14, 41, 6]. We follow the official guidelines to perform zero-shot evaluation on the compared MLLMs. A general language instruction for these methods is as follows:
Suppose you are an <Action_Task>coach. Watch the given video, and answer a paragraph of commentary about what is done well in the execution and what can be improved. Your answer should be like: <Answer_Example>. Pay attention to the execution of <Technical_Dimensions>”,
where <Action_Task> is the action task (e.g., Reverse Layup), <Answer_Example> is an example of the answer (selected from training set), and <Technical_Dimensions> indicates the a list of technical dimensions (e.g., head & eyes, arms & elbows, and human-object interaction) whose corresponded technical keypoints are mentioned in the ground-truth commentary.
B.3 More Qualitative Results

### v2

@@ 5.3 Main Results
Table 5: Ablation studies on the impact of TechPoint Features Augmentation. T.P.Text: TechPoint text features. T.P.Type: TechPoint-type prompts. Quality: quality-related prompts.
Refer to caption Figure 6: Qualitative results on the generated coaching commentary. Our TechCoach generates precise and detailed coaching commentary on what is done well and what can be improved from the given action videos. Correctly matched parts are highlighted in colors.
5.4 Ablation Studies
- Is TechPoint-level alignment necessary? We study this question by gradually ablating the training losses illustrated in Eq. 3 from our full model. The results are shown in Sec. 5.3. By comparing Row 3 with Row 1-2, it can be clearly observed that the model achieves unstable performance when simply combining the $L_{MSE}$ and $L_{MTM}$. However, when further combining the TechPoint-level alignment loss ($L_{Align}$), we observe stable performance improvement on all the compared metrics (Row 4). Results in Fig. 5 further illustrate how $L_{Align}$ impacts the performance. Starting from SwinBERT, after adding all our designs except $L_{Align}$, model performs much better on mentioning the technical details (3.04 improvement on GPT-M Score) due to the TechPoints inputs, but does not obtain similar improvement on perceiving action quality (only 1.45 improvement on GPT-Q Score). After further adding $L_{Align}$, much stronger improvement is observed on GPT-Q Score. All these results indicate the TechPoint-level alignment is necessary and will bring stronger action quality perception ability. - Does TechPoint-level alignment equal to binary classification? The main target of the TechPoint-level alignment (TP-Align) is to ensure the TechPoint-related Quality Embeddings (see Fig. 3 and Sec. 4.4 for more details) carry the quality-related information. To achieve this, a simple solution is to perform binary classifications over the TechPoint-related Quality Embeddings (i.e., classify whether the execution shows strengths or weaknesses on each TechPoint), which we note as TP-CLS. To dive deeper, we conduct a study on the influence of such an alternative solution. As shown in Sec. 5.3, by comparing Row 2 with Row 1, we observe such a solution strengthens the quality perception of the model, resulting in performance improvement on score regression metrics. However, without explicit alignment with the TechPoint-level coaching commentary, Bi-CLS falls short in understanding deeper relationships between the action execution and TechPoints, resulting in a clear performance gap between our TechPoint-level alignment, especially on commentary generation. - Impacts of the TechPoint Query Augmentation. As discussed in Sec. 4.4, we construct TechPoint queries by augmenting the TechPoint text features (T.P.Text) with TechPoint-type prompts (T.P.Type) and quality-related prompts (Quality). In Sec. 5.3, we gradually ablate the components from the augmentation, and the performance deterioration shows the contribution of each component. - More ablation studies. We also provide analysis on other aspects like (i) Impacts of various types of TechPoint-level alignment loss; (ii) Influence of the training videos from different views. Please refer to the Appendix for more details.
5.5 Qualitative Results
To further illustrate the effectiveness of the proposed TechCoach, we provide qualitative results on the generated coaching commentary. As shown in Fig. 6, TechCoach is able to understand various technical TechPoints of different actions and provide detailed commentary on what is done well (e.g., “the extension of the shooting arm and wrist action were also commendable”) and what can be improved (e.g., “improvements are needed in core engagement to prevent being pulled from the wall”) from the given action videos. More qualitative results are shown in the Appendix.
6 Conclusion
We investigate Descriptive Action Coaching, a novel task that aims to provide coaching feedback on what is done well and what can be improved from an action execution. To support this task, we develop an automated pipeline for constructing the EE4D-DescCoach dataset which features clean and detailed coaching commentary on both TechPoint and instance levels. The TechPoint-level commentary provides new supervision for incorporating TechPoint-level reasoning into the action coaching process, encouraging us to build TechCoach, a new framework empowered by a Context-aware TechPoint Reasoner. Besides demonstrating strong performance of the proposed TechCoach, extensive experiments also highlights the effectiveness and the necessity of the proposed Context-aware TechPoint Reasoner. In Appendix, we further discuss the limitations of our work and outline future research directions, including (i) Scaling up TechCoach, (ii) Hierarchical TechPoints, and (iii) Pose-Assisted DescCoach. We expect the proposed new task, dataset, and method provide a new perspective for advancing current AQA to a more explainable and practical scenario.
Appendix A Appendix
In this appendix, we provide more details about the evaluation metrics, implementations, experiment results and analysis in Appendix B. Subsequently, more details about the EE4D-DescCoach dataset are introduced in Appendix C, including examples of general technical points (TechPoints), hierarchical coaching commentary, prompts used for LLM-based annotation collection, and quality ensurance. Finally, we discuss the limitations and potential directions for future work in Appendix D.
Appendix B More Details about Experiments
B.1 More Details about the Evaluation Metrics
Spearman’s rank correlation coefficient ($\rho$) [41] is the most widely used metric to evaluate models in the field of AQA. It measures the correlation between the predicted score series and the ground-truth score series, which is defined as follows: $\rho=\frac{\sum^{N}_{i=1}(x_{i}-\bar{x})(y_{i}-\bar{y})}{\sqrt{\sum^{N}_{i=1}(x_{i}-\bar{x})^{2}\sum^{N}_{i=1}(y_{i}-\bar{y})^{2}}},$ (4) where $x$ and $y$ are the rankings of the two series. Relative-L2 Distance (RL2) [54] evaluates the numerical differences between the predicted and ground-truth scores, which can be written as: $RL2=\frac{1}{N}\sum^{N}_{n=1}(\frac{|s_{n}-\hat{s}_{n}|}{s_{max}-s_{min}})^{2},$ (5) where $s_{n}$ and $\hat{s}_{n}$ indicate the ground-truth and predicted scores, respectively. $s_{max}$ and $s_{min}$ indicate the highest and lowest scores for an action instance. BLEU [28] quantifies the overlap between the generated sentences and the reference sentences by comparing n-grams (contiguous word sequences) shared between them. In this work, we compute the 1-gram-based BLEU score to evaluate the models. METEOR [4] improves upon BLEU by introducing flexible word matching (e.g., synonyms and stemming) to better capture meaning beyond exact matches. Additionally, it adopts a recall-oriented approach to ensure all reference words are considered, and incorporates word order penalties to evaluate fluency and coherence. CIDEr [40] is originally proposed for image captioning task. It measures the similarity between generated and reference texts using n-gram overlap weighted by TF-IDF. This ensures important but less common phrases are emphasized, rewarding consensus among references while penalizing overly generic outputs. BERTScore [61] evaluates natural language generation by comparing the semantic similarity between generated and reference texts using contextual embeddings from pre-trained BERT-like models. Unlike traditional metrics that rely on n-gram overlap, BERTScore computes token-level similarity based on meaning, capturing nuances like synonyms and paraphrasing. In our work, we use the DeBERTa-xlarge [11] to extract token-level embeddings as it shows great correlation with human evaluation. Refer to caption Figure 7: The detailed prompts for LLM-based evaluation. Zoom in for best view. Refer to caption Figure 8: An overview of the Spatial-Temporal Context Enhancer. Here the $\oplus$ indicates the element-level adding operation. GPT-M and GPT-Q are the LLM-based metrics proposed to evaluate whether the generated commentary shares the same technical details and the same coaching opinions as the ground-truth commentary. We employ the GPT-4o-mini to determine the number of both-mentioned technical details and the shared coaching opinions. Detailed prompts are provided in Fig. 7. After that, the GPT-M score is calculated as the averaged proportion of the both-mentioned technical details to those mentioned by ground truth across all testing instances. The GPT-Q score is calculated as the averaged proportion of the technical details sharing the same coaching opinions to the both-mentioned technical details. Refer to caption Figure 9: Qualitative results on the LLM-based evaluation. We take genertaedt and ground-truth commentary from Fig. 5 in the main paper, and prompt LLM to produce accurate counts along with detailed analysis. TD: recognized technical details.
B.2 More Implementation Details
- Spatial-Temporal Context Enhancer. As discussed in Sec. 4.3 of the main paper, inspired by TimeSformer [5], we adopt a Spatial-Temporal Context Enhancer to enhance the spatial-temporal context across the clip-level features. An overview of this module is shown in Fig. 8. Specifically, we use a linear projector to map the clip-level features $F_{clip}$ to the hidden dimension $D$, and feed them into $N_{st}$ Transformer blocks that include spatial- and temporal-attention layers along with the FFN layers. In our implementation, the $N_{st}$ is set to 4. - Other details of TechCaoch. The visual backbone and text encoder are initialized by the pretrained weights and frozen during training. For the visual backbone, we use the pretrained weights of InternVideo2-stage3 [44]. We use the pretrained DeBERTa-xlarge [11] as the text encoder. Following [19, 59], we employ AdamW [21] as the optimizer, and utilize a learning rate warm-up during the early 10% training steps followed by linear decay. $\lambda_{1}$ and $\lambda_{2}$ are set to 0.1 and 0.3, respectively. Due to the characteristics of EgoExo4D [10] and the annotation pipeline of EE4D-DescCoach, it is not possible to ensure each instance includes rounded commentary on every possible TechPoints (e.g., in some cases the coaching commentary would not mention the execution about the head & eyes). Therefore, we make some adjustments on the prompts or instructions for the compared models to guide the models to focus on those TechPoints mentioned in the ground-truth Commentary. For TechCoach, we employ zero-masking on the TechPoint-related quality embeddings corresponded to a specific TechPoint only if both the strength and weakness on this TechPoint are unmentioned in the ground truth. Note that such a design does NOT leak the action quality information, and similar adjustments are applied on the compared models [59, 18, 15, 44, 6] for fair comparisons. - Details about USDL [38]. USDL proposes to convert the score regression problem to a distribution prediction problem. We primarily follow the official implementation. We make adjustments on the output dimension to align with the score range in our experiments. - Details about CoRe [54] and TPT [3]. These two methods adopt the Contrastive Regression paradigm to generate action quality scores by predicting the score differences between the query instance and multiple exemplar instances. Our experiments are conducted based on the official implementations of these methods. The exemplars are randomly selected from the training instances with the same action task (e.g., Reverse Layup) as the query instance. The number of selected exemplars is set to 10. - Details about SwinBERT [19]. SwinBERT proposes a general framework for video captioning. In addition to the modification of the visual backbone, we also modified the length of masking tokens during training to ensure consistency with our TechCoach. Additionally, following [59], we add an extra full-connected layer to regress the score. - Details about PGMI [59]. PGMI is a prompt-guided framework for Multi-Task Learning AQA paradigm. Our experiments on PGMI are conducted using the official implementation. We adjust the number of score groups to 10 as the score range in our experiments is in [0, 10]. Besides, we modify the input templates for the text generator to include the score information and the technical dimensions that should be focused on. Specifically, the template is adjusted as follows: The player is practicing <Action_Task>, and gets a score of <Predicted_Score>. Pay attention to the execution of <Technical_Dimensions>”, where <Action_Task> is the action task, <Predicted_Score> indicates the predicted score by the model, and <Technical_Dimensions> indicates the a list of technical dimensions (e.g., head & eyes, arms & elbows, and human-object interaction) whose corresponded TechPoints are mentioned in the ground truth. - Details about the Compared MLLMs [18, 15, 44, 6]. We follow the official guidelines to perform zero-shot evaluation on the compared MLLMs. A general language instruction for these methods is as follows: Suppose you are an <Action_Task>coach. Watch the given video, and answer a paragraph of commentary about what is done well in the execution and what can be improved. Your answer should be like: <Answer_Example>. Pay attention to the execution of <Technical_Dimensions>”, where <Action_Task> is the action task (e.g., Reverse Layup), <Answer_Example> is an example of the answer (selected from training set), and <Technical_Dimensions> indicates the a list of technical dimensions (e.g., head & eyes, arms & elbows, and human-object interaction) whose corresponded TechPoints are mentioned in the ground-truth commentary. In Sec. B.3, we also explore building a DescCoach-specific MLLM. Specifically, we take VideoChat2 [15] as an example, and fine-tune it with LoRA [12] while keep the other parameters frozen on our EE4D-DescCoach dataset. Refer to caption Figure 10: User study. Our TechCoach wins the compared baselines frequently in human evaluation. Refer to caption Figure 11: Qualitative results on the LLM-based evaluation. We take genertaedt and ground-truth commentary from Fig. 5 in the main paper, and prompt LLM to produce accurate counts along with detailed analysis
{NiceTabular} l|cccccc Methods Commentary
[…]
VideoChat2-7B [15] 23.18 2.98 11.59 61.37 14.09 22.28
VideoChat2-7B-FT [15] 28.74 4.88 13.27 62.98 36.32 23.62
TechCoach(Ours) 37.06 14.62 16.39 66.89 50.44 38.15
Table 6: Comparison with fine-tuned MLLM. While finetuning VideoChat2 on our EE4D-DescCoach dataset bring stable improvement on all metrics, the modest gain on GPT-Q score indicates that the model still struggle to obtain strong action quality perceiving ability. FT: finetuned on our EE4D-DescCoach dataset.
{NiceTabular} l|cc|cccc Variants Score   Commentary
[…]
InfoNCE-Align 71.38 4.60 36.76 13.72 16.26 65.78
MarginRanking-Align 71.67 4.36 36.96 14.29 16.35 66.45
L2-Align 72.46 4.35 37.06 14.62 16.39 66.89
Table 7: Comparison with variants with other types of feature alignment losses. TechCoach achieves stronger performance with simple yet effective L2-distance-based alignment loss.
{NiceTabular} l|cc|cccc Variants Score   Commentary
@@ B.2 More Implementation Details
B.3 More Comparison Results and Analysis
- User study. To further compare the performance across different models, we conducted a comprehensive user study comparing our TechCoach against three baselines (i.e., PGMI[59], SwinBERT[19], and InternVL2-8B[6]). We randomly select 50 instances from evaluation set and ask all four models to generate parallel outputs. We recruit five labelers (not involved in our project) and each of them is assigned instances through random sampling to eliminate evaluation bias. We provide similar assessment instructions as those for LLM evaluation (see Fig. 7), and ask the labelers performed blind assessments by ranking the four outputs per instance based on semantic alignment with ground-truth (GT) text, allowing tied rankings (e.g., TechCoach=PGMI>InternVL2-8B>SwinBERT). After that, we computed pairwise comparison metrics: (1) Win Rate: Percentage of instances where TechCoach outperforms the baseline; (2) Tie Rate: Percentage of instances where there is a tie between TechCoach and the baseline; (3) Loss Rate: Percentage of instances where the baseline model prevail. The user study results are shown in Fig. 10. Compared with other baselines, our TechCoach achieves clearly higher Win Rate, further illustrating the generated text of TechCoach can effectively catch the semantic meaning mentioned in Groun-Truth and align with human preference. - Adding general TechPoints into the prompts for MLLMs. As discussed in Sec. 4.2, our TechCoach inferences with assumption that general TechPoints (GTP) are available. To evaluate the influence of GTP for general MLLMs, we try to add general TechPoints into the prompts for MLLMs. As shown in Fig. 11, adding the (GTP) into the prompts brings unconsistent performance effects across different MLLMs (i.e., InternVL2-8B gets improvement but VideoChat2-7B and InternVideo2 gets opposite effects). These results suggest that a more effective approach to integrating TechPoints into the action assessment process is necessary. - Fine-tuning MLLM on EE4D-DescCoach dataset. As another branch of exploration, we try to fine-tune MLLM on our EE4D-DescCoach dataset. We take VideoChat2 [15] as an example, and fine-tune it with LoRA [12]. The results are shown in Sec. B.2. While finetuning VideoChat2 on our EE4D-DescCoach dataset brings stable improvement on all metrics, the modest gain on GPT-Q score indicates that the model still struggle to obtain strong action quality perceiving ability. - Why does InternVL2 achieve abnormal performance on traditional metrics? In Tab. 2 of main paper, InternVL2 [6] family models obtain abnormal performance on traditional n-gram based metrics (i.e., BLEU, METEOR, CIDEr) and feature similarity-based metric (BERTScore). By going deeper to the outputs of InternVL2, we find the reasons behind it could be: even if an answer sample is provided in the prompt, InternVL2 tends to provide answers in a fixed, list-like format. For example: ## Execution Done Well: 1. **Starting Position**: The player demonstrates a strong starting position with … 2. **Ball Handling**: The dribbling technique appears consistent … The formats of answers and ground-truth commenatry are highly inconsistent, which makes the traditional metrics fail to effectively evaluate the model performance.
B.4 More Ablation Studies
@@
- Impacts of various types of TechPoint-level alignment loss. As discussed in Sec. 4.4, L2 distance is used to measure the distance between TechPoint-related quality embeddings and TechPoint-level commentary features. We further investigate the imapcts of alignment loss by introducing two other kinds of alignment losses, i.e., InfoNCE and MarginRanking. As shown in LABEL:{tab:ablation_alignment_loss}, TechCoach achieves better performance with the L2 distance-based alignment loss. - Impacts of videos from various views. As discussed in Sec. 5.2, we train TechCoach with both Egocentric and Exocentric videos. Here we conduct an experiment on the influence of the training videos from different views. As shown in Sec. B.2, Ego-Only and Exo-Only variants showcase comparable performance. By combining videos from different views, TechCoach achieves stronger performance, which shows that egocentric and exocentric viewpoints are complementary for action coaching.
B.5 More Qualitative Results